OpenStudy (anonymous):

sin(2cos^-1(-3/5)) is

OpenStudy (anonymous):

let 2cos(inv) (-3/5) = A cos A/2=(-3/5) sinA=2sinA/2 * cosA/2 = 2sinA/2 * (-3/5) =(-6/5) * sqrt(1-cos^2 (A/2)) = (-6/5)*(4/5) = -24/25 is it correct???

OpenStudy (anonymous):

@Sarkar - thanks...you got it right.:D

OpenStudy (mertsj):

\[\sin2x=2sinxcosx=2(\frac{4}{5})(\frac{-3}{5})=\frac{-24}{25}\]

OpenStudy (anonymous):

But did you get it??

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

good...